• Compute the number of modes that will be 

supported by a fibre with core diameter of 20mm, a 

core refractive index of 1.5 and a cladding index of 

1.48 if it is excited by light of frequency 0.25MHz


1
Expert's answer
2021-03-05T07:25:28-0500

Answer

Given radius of core

a="D\/2" =0.01m

"n_1=1.5\\n_2=1.48"

Frequency f="0.25\times10^6Hz"

So wavelegth

"\lambda=\frac{c}{f}=1200m"

Therefore no. Of modes

"V=\frac{2\times3.14\times0.01}{1200}\sqrt{(1.5)^2-(1.48)^2}"

"=1.277\times10^{-5}"

According concepts

V"<2.405"

Therefore only one mode is possible.